### v1

@@ Abstract
Representations from large language models are highly effective at predicting BOLD fMRI responses to language stimuli. However, these representations are largely opaque: it is unclear what features of the language stimulus drive the response in each brain area. We present generative explanation-mediated validation (GEM-V), a framework for generating concise explanations of language selectivity in the brain and then validating those explanations in follow-up experiments that use synthetic stimuli. This approach is successful at explaining selectivity both in individual voxels and cortical regions of interest (ROIs). We show that explanatory accuracy is closely related to the predictive power and stability of the underlying statistical models. These results demonstrate that LLMs can be used to bridge the widening gap between data-driven models and formal scientific theories.
@@ 1   Introduction
To overcome this challenge, we introduce the generative explanation-mediated validation (GEM-V) framework. GEM-V translates deep learning models of language selectivity in the brain into concise verbal explanations, and then designs follow-up experiments to verify that these explanations are causally related to brain activity.
Figure 1: Driving single-voxel response with generative explanation-mediated validation. (a) Voxelwise BOLD responses were recorded using fMRI as human subjects listened to 20 hours of narrative stories. An encoding model $f$ was fit to predict these responses from the story text. $f$ consists of a linear model fit on representations extracted from an LLM, which are not readily interpretable. Encoding models were tested by predicting responses on held-out fMRI data, and only well-performing models were selected for further analysis. (b) We used an automated procedure to find a verbal description of the function that $f$ computes for each voxel. First, we tested $f$ on a large catalog of $n$-grams ($n=1,2,3)$ and found those that maximally drove predicted responses. These $n$-grams were then summarized into stable explanation candidates using a powerful instruction-tuned LLM. Finally, we evaluated each explanation candidate by generating corresponding synthetic sentences and testing that these sentences yielded large predictions from $f$. (c) To test whether the generated explanations were causally related to activation in the brain, we used an LLM to produce synthetic narrative stories where each paragraph is designed to drive responses based on the generated explanation for one voxel. For each subject we constructed stories to drive 17 well-modeled voxels with diverse selectivity. These stories were then presented to the subjects in a second fMRI experiment. (d) Average BOLD response during its driving paragraph for each voxel, relative to baseline, i.e. average response to all non-driving paragraphs. On average, driven responses were significantly higher than baseline for each subject ($p=0.020$ (S01), $p<10^{-5}$ (S02), $p=0.009$ (S03); permutation test, FDR-corrected). For well-driven voxels, this means that the generated explanation is causally related to activation of that voxel, and thus that we have successfully translated the LLM-based encoding model into a verbal explanation. (e) Average BOLD response for each selected voxel to each of the driving paragraphs in one subject (S02). Responses to the driving paragraph generated using the explanation for that voxel appear along the main diagonal. Explanations that were used to construct the driving paragraphs are shown below. BOLD responses were generally high for the driving paragraphs for each voxel as well as semantically related paragraphs (e.g. directions and locations, emotional expression and laughter).
@@ 2   Language encoding models
We designed GEM-V to interpret language encoding models, which are data-driven, LLM-based models of cortical language selectivity. We thus began by fitting encoding models (Fig. 1a) for each voxel in each of 3 human subjects using 20 hours of passive language listening fMRI data per subject collected in a prior study [11]. The 20 hours of speech stimuli were transcribed and then passed through multiple open-source large language models [12, 13] to obtain activation vectors for each word. Multiple models were built using different LLMs to ensure that predictions are stable [14, 15], i.e. not idiosyncratic to a single model. Regularized linear regression was then used to predict the response timecourse of each voxel as a weighted combination of LLM activations. Encoding models were tested by predicting responses on held out fMRI data and then computing the correlation between predicted and actual responses. Encoding models using advanced LLMs achieve very high prediction performance [8]; here, only models with sufficient prediction performance ($r>0.15$) were used in further analyses, with many voxels predicted at $r>0.7$. Full preprocessing, modeling, and data collection details are provided in the Methods.
3   Generating and evaluating natural-language explanations for single voxels
@@
Each encoding model represents the language selectivity of one part of the brain, but does so as a linear combination of LLM activations that are not human-interpretable. The first step in GEM-V is to convert these models into concise natural language explanations: a word or short phrase summarizing the properties of language that drive responses most strongly. This was accomplished by using an instruction-finetuned LLM [16] to generate candidate explanations by summarizing the n-grams that yield the largest predictions from the model (Fig. 1b) [17]. The resulting explanation is given in easily understandable natural language. For example, we found that some voxels appeared to be selective for language describing Food preparation (Fig. 1b). To ensure that the generated explanation correctly captures the function computed by the encoding model, we used the same LLM to generate synthetic text from the explanation and checked that this text could successfully drive the voxel encoding model. However, this does not guarantee that this function is correctly aligned to the activity of the brain in the real world.
@@
To fully close the loop and confirm that a generated explanation accurately reflects language selectivity in the brain, the second step of GEM-V is to automatically design a new neuroimaging experiment to test the generated explanation in vivo. This was done by prompting an instruction-finetuned LLM [16] to generate narratives that should selectively drive cortical activation based on that explanation (Fig. 1c). If the explanation generated for a voxel is accurate, then text generated according to that explanation should elicit large responses in that voxel; this would demonstrate that the explanation has causal influence over that voxel’s activity. Given a selected set of voxels and their explanations, we constructed narrative stories by iteratively prompting an LLM [16] to prioritize a different voxel’s explanation as the main focus of each paragraph. This allowed us to test whether voxels selectively respond to paragraphs that match their explanation. The LLM ensures these stories remain coherent and engaging, helping to keep subjects attentive during the fMRI experiment [18, 7].
@@
We used GEM-V to generate and test explanations for 17 voxels with strong encoding model predictive performance in each of 3 subjects. The number of voxels was selected so that resulting narratives would be similar in length to the stimuli used to initially fit the encoding models [11]. We then measured the average response of each target voxel during the paragraph that was designed to drive it and compared this to the average response to other paragraphs (Fig. 1d). Across voxels, responses to driving paragraphs were significantly greater than baseline responses for each subject ($p=0.020$ (S01), $p<10^{-5}$ (S02), $p=0.009$ (S03); permutation test with Benjamini-Hochberg false discovery rate correction [19]). Of the 51 tested voxels across the 3 subjects, 41 show increased response, with an average increase of 0.198 standard deviations over baseline. Differences in effect size are related to head motion of subjects during test time and variation in sample size (Table 2). This demonstrates that the generated explanations are mostly effective drivers of brain activity for their chosen voxels. Further, because this experiment was not based directly on the LLM encoding model [9] but was designed using the generated explanation, this suggests that we have successfully found explanations that are causally linked to brain activity.
@@
Figure 2: Driving ROI response with generative explanation-mediated validation for subject S02. (a) Explanations were generated and used to drive 8 well-defined regions of interest. Responses in all ROIs were significantly driven above baseline ($p<0.05$; permutation test, FDR-corrected). (b) To understand driving performance with more granularity, we color each voxel in each ROI by how well it was driven by its corresponding driving paragraph. The resulting composite flatmap occasionally shows subregions within ROIs that are more selectively driven for a particular explanation. (c) GEM-V can also be used to build more nuanced theories of cortical semantic selectivity. We focused on three ROIs that are known to have similar selectivity for place concepts: retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). When explanations were generated for each ROI independently we found that each ROI was driven by all three driving paragraphs (left side). To distinguish these ROIs, we used GEM-V to find new explanations and construct stories that would selectively drive each area while suppressing the other two. Testing these stories in an fMRI experiment showed that we succeeded in finding selective explanations for two ROIs: RSC is selectively driven by location names and PPA by unappetizing foods. However, the explanation for OPA, spatial positioning & directions, drove responses in all three ROIs (right side). (d) Visualization of the place area driving experiment with voxel-level granularity. We show a 3-channel flatmap showing the outcome of each location-selective driving experiment; a voxel is more red/green/blue if that voxel was driven by the corresponding ROI explanation.
@@ 4   Explaining selectivity in regions of interest (ROIs)
Theories of language selectivity often revolve around regions of interest (ROIs) that are believed to be selective for specific semantic categories [20], such as locations (OPA) [21, 22] or body parts (EBA) [23, 24]. We explored whether GEM-V could be used to independently uncover the semantic selectivity of regions of interest, both for generating new hypotheses as well as re-establishing known functional selectivity. We used GEM-V to find explanations and design synthetic driving stimuli for a diverse selection of individual ROIs that were localized using separate stimuli: the extrastriate body area (EBA), intraparietal sulcus (IPS), occipital face area (OFA), posterior superior temporal sulcus (pSTS), superior premotor ventral (sPMv), retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). These ROIs were identified in each subject using separate localizer scans. Fig. 2a shows the generated explanations and average response above baseline for each ROI during each driving experiment. The explanations found by GEM-V and validated in the follow-up experiment are well-matched to known selectivity, e.g. Body parts in EBA, and Scenes and settings in PPA. Driving succeeded for every ROI (all $p<0.05$; permutation test with FDR correction), although for some ROIs, e.g. RSC, the effect is stronger. Fig. 2b visualizes driving responses for each ROI on a composite flatmap, where each voxel is colored on the basis of how much it was driven by its ROI’s explanation. This map suggests that the variability in driving between ROIs reflects functional heterogeneity within each region.
@@
We next tested where GEM-V could be used to describe differences between regions that appear to have similar selectivity [20]. Examining three ROIs known to be selective for places or locations (RSC, PPA, OPA) [25], we found that the driving paragraphs designed independently for each ROI also drove responses in the others (Fig. 2c left). To differentiate these ROIs, we used GEM-V to generate new explanations that should selectively drive each one of these regions while suppressing the other two [26]. These selective explanations deviated somewhat from the original ROI explanations, e.g. for RSC the explanation changed from Travel and location names to just Location names, suggesting that the travel part of the explanation is common across ROIs while location names are more unique to RSC. New stories were generated to include only instances of the corresponding explanation, for instance, when driving only RSC, location names were included, while directions were specifically excluded (see prompts in Section A.1). Since these two categories commonly cooccur in natural language, these synthetic stimuli enabled us to examine differences that might not be present in typical naturalistic studies. Testing these stories in another fMRI experiment showed that GEM-V was able to find explanations that can selectively drive two of the ROIs—Location names for RSC and Unappetizing foods for PPA—while the explanation of Spatial positioning & directions for OPA still drove responses in all three ROIs (Fig. 2c right; Fig. 2d). These results demonstrate that GEM-V can be used to build more nuanced theories of cortical semantic selectivity beyond general trends that are well-understood.
Figure 3: Responses to driving paragraphs reproduce known semantic contrasts and open new hypotheses. To understand how each explanation activates the entire cortex, we visualize the driven response (average normalized BOLD response) for six explanations. (a) Some maps recovered well-established results, e.g. the explanation Locations gave strong responses in the place areas RSC, OPA, and PPA. (b) Other maps independently confirmed newer hypotheses, e.g. the explanation Food Preparation activates a region in ventral occipital cortex near the fusiform face area (FFA) [27, 28]. (c) Finally, some explanations did not clearly map to any known hypotheses, but may suggest new directions for future research. All observed activation patterns along with ROI localization explanations are presented in Appendix A.7.
5   Responses to synthetic stories reveal underlying contrast effects
Although the stories generated by GEM-V are designed to activate single voxels or ROIs, we can measure how much each explanation activates every part of cortex. These experiments can thus be seen as a hybrid between classical block-designs that are easy to analyze and natural stimulus experiments that elicit stronger and more widespread activation in cortex [18, 7]. To demonstrate this, we computed the average activation map for each explanation by averaging responses over timepoints within each paragraph. The results for six selected explanations in one subject are shown in Fig. 3.
Some explanations were similar to previously well-established contrasts (Fig. 3a) and gave the expected results. For example, the explanation Locations gave strong responses in the place areas RSC, OPA, and PPA [25]. Others confirmed newer hypotheses. For example, the Food Preparation explanation (Fig. 3b Top) activates a region in ventral occipital cortex near the fusiform face area (FFA). Selectivity for this area to food was found recently in responses to images [27, 28], but these results demonstrate for the first time that this selectivity also extends to linguistic descriptions of food. Finally, some explanations did not clearly map to any known hypotheses (Fig. 3c), but still demonstrate suggestive patterns of activation that could be explored in future experiments. Maps for other explanations are provided in Appendix A.7.
6   Factors underlying effective GEM-V stories and voxels

### v2

@@ Abstract
Representations from large language models are highly effective at predicting BOLD fMRI responses to language stimuli. However, these representations are largely opaque: it is unclear what features of the language stimulus drive the response in each brain area. We present generative causal testing (GCT), a framework for generating concise explanations of language selectivity in the brain from predictive models and then testing those explanations in follow-up experiments using LLM-generated stimuli. This approach is successful at explaining selectivity both in individual voxels and cortical regions of interest (ROIs), including newly identified microROIs in prefrontal cortex. We show that explanatory accuracy is closely related to the predictive power and stability of the underlying predictive models. Finally, we show that GCT can dissect fine-grained differences between brain areas with similar functional selectivity. These results demonstrate that LLMs can be used to bridge the widening gap between data-driven models and formal scientific theories.
@@ 1   Introduction
To overcome this challenge, we introduce the generative causal testing (GCT) framework. GCT translates predictive deep learning models of language selectivity in the brain into concise verbal explanations, and then designs follow-up experiments using LLM-generated stimuli to verify that these explanations are causally related to brain activity.
Figure 1: Driving single-voxel response with generative causal testing. (a) Voxelwise BOLD responses were recorded using fMRI as human subjects listened to 20 hours of narrative stories. An encoding model $f$ was fit to predict these responses from the story text. $f$ consists of a linear model fit on representations extracted from an LLM, which are not readily interpretable. Encoding models were tested by predicting responses on held-out fMRI data, and only well-performing models were selected for further analysis. (b) We used an automated procedure to find a verbal description of the function that $f$ computes for each voxel. First, we tested $f$ on a large catalog of $n$-grams ($n=1,2,3)$ and found those that maximally drove predicted responses. These $n$-grams were then summarized into stable explanation candidates using a powerful instruction-tuned LLM. Finally, we evaluated each explanation candidate by generating corresponding synthetic sentences and testing that these sentences yielded large predictions from $f$. (c) To test whether the generated explanations were causally related to activation in the brain, we used an LLM to produce synthetic narrative stories where each paragraph is designed to drive responses based on the generated explanation for one voxel. For each subject we constructed stories to drive 17 well-modeled voxels with diverse selectivity. These stories were then presented to the subjects in a second fMRI experiment. (d) Average BOLD response during its driving paragraph for each voxel, relative to baseline, i.e. average response to all non-driving paragraphs. On average, driven responses were significantly higher than baseline for each subject ($p=0.020$ (S01), $p<10^{-5}$ (S02), $p=0.009$ (S03); permutation test, FDR-corrected). For well-driven voxels, this means that the generated explanation is causally related to activation of that voxel, and thus that we have successfully translated the LLM-based encoding model into a verbal explanation. (e) Average BOLD response for each selected voxel to each of the driving paragraphs in one subject (S02). Responses to the driving paragraph generated using the explanation for that voxel appear along the main diagonal. Explanations that were used to construct the driving paragraphs are shown below. BOLD responses were generally high for the driving paragraphs for each voxel as well as semantically related paragraphs (e.g. directions and locations, emotional expression and laughter).
@@ 2   Language encoding models
We designed GCT to interpret language encoding models, which are data-driven, LLM-based models of cortical language selectivity. We thus began by fitting encoding models (Fig. 1a) for each voxel in each of 3 human subjects using 20 hours of passive language listening fMRI data [11]. The 20 hours of speech stimuli were transcribed and then passed through open-source large language models [12, 13] to obtain activation vectors for each word. Multiple models were built using different LLMs to ensure that predictions are stable [14, 15], i.e. not idiosyncratic to a single model. Regularized linear regression was then used to predict the response timecourse of each voxel as a weighted combination of LLM activations. Encoding models were tested by predicting responses on held out fMRI data and then computing the correlation between predicted and actual responses. Encoding models using advanced LLMs achieve very high prediction performance [8]; here, only models with sufficient prediction performance ($r>0.15$) were used in further analyses, with many voxels predicted at $r>0.7$. Full preprocessing, modeling, and data collection details are provided in the Methods.
3   Generating and testing natural-language explanations for single voxels
@@
Each encoding model represents the language selectivity of one part of the brain, but does so as a linear combination of LLM activations that are not human-interpretable. The first step in GCT is to convert these models into concise natural language explanations: a word or short phrase summarizing the properties of language that drive responses most strongly. This was accomplished by using an instruction-finetuned LLM [16] to generate candidate explanations by summarizing the n-grams that yield the largest predictions from the model (Fig. 1b) [17]. The resulting explanation is given in easily understandable natural language. For example, we found that some voxels appeared to be selective for language describing Food preparation (Fig. 1b). To ensure that the generated explanation correctly captures the function computed by the encoding model, we used the same LLM to generate synthetic text from the explanation and checked that this text could successfully drive the voxel encoding model. However, this does not guarantee that this function is correctly aligned to the activity of the brain in the real world.
@@
To fully close the loop and confirm that a generated explanation accurately reflects selectivity in the brain, the second step of GCT is to automatically design a new neuroimaging experiment to test the generated explanation in vivo. This was done by prompting an instruction-finetuned LLM [16] to generate narratives that should selectively drive cortical activation based on that explanation (Fig. 1c). If the explanation generated for a voxel is accurate, then text generated according to that explanation should elicit large responses in that voxel; this would demonstrate that the explanation has causal influence over that voxel’s activity. Given a selected set of voxels and their explanations, we constructed narrative stories by iteratively prompting an LLM [16] to prioritize a different voxel’s explanation as the main focus of each paragraph. This allowed us to test whether voxels selectively respond to paragraphs that match their explanation. The LLM ensures these stories remain coherent and engaging, helping to keep subjects attentive during the fMRI experiment [18, 7].
@@
We used GCT to generate and test explanations for 17 voxels with strong encoding model predictive performance in each of 3 subjects. The number of voxels was selected so that resulting narratives would be similar in length to the stimuli used to initially fit the encoding models [11]. We then measured the average response of each target voxel during the paragraph that was designed to drive it and compared this to the average response to other paragraphs (Fig. 1d). Across voxels, responses to driving paragraphs were significantly greater than baseline responses for each subject ($p=0.020$ (S01), $p<10^{-5}$ (S02), $p=0.009$ (S03); permutation test with Benjamini-Hochberg false discovery rate correction [19]). Of the 51 tested voxels across the 3 subjects, 41 show increased response, with an average increase of 0.198 standard deviations over baseline. Differences in effect size are related to head motion of subjects during test time and variation in sample size (Table 2). This demonstrates that the generated explanations are mostly effective drivers of brain activity for their chosen voxels. Further, because this experiment was not based directly on the LLM encoding model [9] but was designed using the generated explanation, this suggests that we have successfully found explanations that are causally linked to brain activity.
@@
Figure 2: Driving ROI response with generative explanation-mediated validation for subject S02. (a) Explanations were generated and used to drive 8 well-defined regions of interest. Responses in all ROIs were significantly driven above baseline ($p<0.05$; permutation test, FDR-corrected). (b) To understand driving performance with more granularity, we color each voxel in each ROI by how well it was driven by its corresponding driving paragraph. The resulting composite flatmap occasionally shows subregions within ROIs that are more selectively driven for a particular explanation. (c) GCT can also be used to build more nuanced theories of cortical semantic selectivity. We focused on three ROIs that are known to have similar selectivity for place concepts: retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). When explanations were generated for each ROI independently we found that each ROI was driven by all three driving paragraphs (left side). To distinguish these ROIs, we used GCT to find new explanations and construct stories that would selectively drive each area while suppressing the other two. Testing these stories in an fMRI experiment showed that we succeeded in finding selective explanations for two ROIs: RSC is selectively driven by location names and PPA by unappetizing foods. However, the explanation for OPA, spatial positioning & directions, drove responses in all three ROIs (right side). (d) Visualization of the place area driving experiment with voxel-level granularity. We show a 3-channel flatmap showing the outcome of each location-selective driving experiment; a voxel is more red/green/blue if that voxel was driven by the corresponding ROI explanation.
@@ 4   Explaining selectivity in regions of interest (ROIs)
Theories of language selectivity often revolve around regions of interest (ROIs) that are believed to be selective for specific semantic categories [20], such as locations (OPA) [21, 22] or body parts (EBA) [23, 24]. We explored whether GCT could be used to independently uncover the semantic selectivity of regions of interest, both for generating new hypotheses as well as re-establishing known functional selectivity. We used GCT to find explanations and design synthetic driving stimuli for a diverse selection of individual ROIs that were localized using separate stimuli: the extrastriate body area (EBA), intraparietal sulcus (IPS), occipital face area (OFA), posterior superior temporal sulcus (pSTS), superior premotor ventral (sPMv), retrosplenial cortex (RSC), the parahippocampal place area (PPA), and the occipital place area (OPA). These ROIs were identified in each subject using separate localizer scans. Fig. 2a shows the generated explanations and average response above baseline for each ROI during each driving experiment. The explanations found by GCT and validated in the follow-up experiment are well-matched to known selectivity, e.g. Body parts in EBA, and Scenes and settings in PPA. Driving succeeded for every ROI (all $p<0.05$; permutation test with FDR correction), although for some ROIs, e.g. RSC, the effect is stronger. Fig. 2b visualizes driving responses for each ROI on a composite flatmap, where each voxel is colored on the basis of how much it was driven by its ROI’s explanation. This map suggests that the variability in driving between ROIs reflects functional heterogeneity within each region.
@@
We next tested where GCT could be used to describe differences between regions that appear to have similar selectivity [20]. Examining three ROIs known to be selective for places or locations (RSC, PPA, OPA) [25], we found that the driving paragraphs designed independently for each ROI also drove responses in the others (Fig. 2c left). To differentiate these ROIs, we used GCT to generate new explanations that should selectively drive each one of these regions while suppressing the other two [26]. These selective explanations deviated somewhat from the original ROI explanations, e.g. for RSC the explanation changed from Travel and location names to just Location names, suggesting that the travel part of the explanation is common across ROIs while location names are more unique to RSC. New stories were generated to include only instances of the corresponding explanation, for instance, when driving only RSC, location names were included, while directions were specifically excluded (see prompts in Section A.1). Since these two categories commonly cooccur in natural language, these synthetic stimuli enabled us to examine differences that might not be present in typical naturalistic studies. Testing these stories in another fMRI experiment showed that GCT was able to find explanations that can selectively drive two of the ROIs—Location names for RSC and Unappetizing foods for PPA—while the explanation of Spatial positioning & directions for OPA still drove responses in all three ROIs (Fig. 2c right; Fig. 2d). These results demonstrate that GCT can be used to build more nuanced theories of cortical semantic selectivity beyond general trends that are well-understood.
5   Discovering new functional ROIs using GCT
The previous results showed that GCT can evaluate explanations for known ROIs, so we next asked if GCT can also discover new functional ROIs. To take advantage of GCT’s strengths, we focused our search on very small ‘micro ROIs’ in prefrontal cortex. Micro ROIs are particularly easy to miss in group fMRI analyses [27], and earlier work has suggested that prefrontal cortex may contain many small specialized functional areas  [20, 28]. Discovering such regions would help develop a more granular picture of cortical functional organization, but this discovery process can be quite challenging owing to the expansive hypothesis space of potential explanations.
We generated candidate ROIs across PFC by selecting small circular regions of cortex for two subjects. This resulted in 582 candidate ROIs for S02 and 535 candidate ROIs for S03. We then ran the GCT explanation pipeline (Fig. 1a-b) to generate explanations for each candidate ROI. We filtered out explanations with a stability score below 0.6 (see Section 9); this resulted in 73 stable candidate ROIs for S02 and 128 stable candidate ROIs for S03. Finally, we grouped the explanations into 4 common categories using keyword search and built GCT stories to evaluate the driving scores of those explanations. These stories were used in a new fMRI experiment.
We found that many of the candidate ROIs were driven as expected (Fig. 3a). Among the selected candidate ROIs, 47 were significantly driven for S02 and 21 were significantly driven for S03 ($p<0.05$, 1-tailed t-test with FDR correction).
Fig. 3b shows the driving scores for a selection of these candidate ROIs across the two subjects. Many of the candidate ROIs are in consistent locations across subjects. The first set of candidate ROIs are selective for the concept of “recognition” and are found in a small region near Brodmann area 9 in medial prefrontal cortex. The selectivity is specific to instances where one person is described to have recognized or noticed another (e.g. my friends saw me, the guard spotted us). This new finding may relate to earlier findings linking this broad region to theory of mind [29]. However, our results show an unexpectedly high degree of specificity both anatomically and semantically. The second set of candidate ROIs are found near the right hemisphere analogue of Broca’s area and are selective for words that signal dialogue between two parties (such as said or told). The proximity of this ROI, which is selective for language describing speech acts, to Broca’s area, which carries out speech acts, enhances grounding theories of cortical organization [30]. The third set of candidate ROIs are near Brodmann area 46 and are selective for mentions of times (e.g. one o’clock). This, along with similar selectivity found in the final set of candidate ROIs for numeric measurement words (e.g. fifty feet) refines research that has found this area may be involved in numeric processing or calculation [31]. The similarities between these two selectivities extend past this single localized region and suggest the presence of a network of regions in right PFC that is specialized for numeric processing. Taken together, these new micro-ROIs provide a refined understanding of semantic processing across PFC.
6   Adversarially testing functional similarity with GCT
Figure 3: Evaluating hypothesized micro-ROIs in prefrontal cortex using GCT. (a) To measure GCT’s ability to aid in the discovery of new brain regions, spatially-contiguous candidate ROIs were defined in a grid pattern throughout prefrontal cortex. Candidate ROIs with high stability scores (see Section 7) were filtered out to define stable ROIs. GCT was used to automatically generate explanations and driving stimuli for these ROIs. The ROI responses to the driving stimuli were then measured in an fMRI experiment, and their average driving scores are shown in a histogram for subject S02 and S03. In both subjects, high-stability candidate ROIs are driven using their corresponding explanation at statistically significant rates; 47 out of 73 are significantly driven for S02 and 21 out of 128 are significantly driven for S03 ($p<0.05$, 1-tailed t-test, FDR-corrected). (b) Significantly driven candidate ROIs across the two subjects are visualized, colored by their driving scores. Significant candidate ROIs for various explanations (Recognition, Time, Dialogue, and Measurements) are in consistent locations across the subjects, suggesting population level trends. (c) We next examined whether GCT can validate functional similarity claims between 5 regions in the language network, localized using an established localizer [32]. We concatenated the driving scores of each of these ROIs into a vector and then computed the cosine similarity between the vectors for different ROIs. We find that the driving vectors of language network ROIs are on average more similar to each other than over 94% of randomly selected pairs of similarly-sized ROIs in UTS02 and over 99% of randomly selected pairs in UTS03 ($p<10^{-5}$; 1-tailed $t$-test), supporting claims of functional similarity across the language network.
Proving that two brain areas have different functional selectivity requires only that the areas differ along one functional dimension. But proving that two brain areas have the same selectivity is much more difficult, as one must show that there is no dimension that differentiates the areas. GCT offers a potential solution to this issue, because it can adversarially search for stimuli that differentially activate brain areas as we showed in Fig. 2. Provided that we have high confidence that we would be able to generate such a stimulus if it existed, then our failure to do so would be affirmative proof of the similarity of the two regions.
As a demonstration of this use case, we focused on the putative ‘language network’ [33], a set of regions that respond readily to language and are claimed to be functionally indistinguishable. First, we localized five language regions in each of two subjects using established methods [32]. For each region, we then used GCT to generate stimuli that would drive that region while minimizing the response in the other regions, as we did successfully in Fig. 2c. Using this approach, we found no stimuli that could drive one region while suppressing the average of the others, strongly supporting claims of functional homogeneity in the language network. Fig. 3c summarizes our results. We found that the driving behaviors of language network ROIs are more similar to each other, as measured by cosine similarity across explanations, than over 94% of randomly selected pairs of similarly-sized ROIs in UTS02 and 99% of all pairs in UTS03 ($p<10^{-5}$; 1-tailed t-test). This result both serves to validate claims of functional similarity across the language network [33] and to demonstrate that GCT can be used as a powerful tool for detecting functional similarity as well as functional differences across ROIs. A full dissection of these language network driving experiments is given in Appendix A.8.
7   Factors affecting GCT’s ability to drive a voxel